40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 301 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  1743   Tue Jul 14 14:54:19 2009 steveConfigurationComputersfb40m2 in 1Y6

Alex and Steve,

SunFire x4600 ( not  MEGATRON 2 , it is fb40m2 ) and JetStor ( 16 x 1 TB drives ) were installed on side rails at the bottom of 1Y6

We cleaned up the fibres and cabling in 1Y7 also

  1742   Tue Jul 14 00:57:11 2009 AlbertoUpdateLockingphotodiode alignment check

Since lately the alignment of the input beam to the interferometer has changed, I went checking the alignment of the beam on the photodiodea. They were all fine except for pd9, that is AS DD 199. Here the DC is totally null. The beam seems to go right on the diode but the scope on the PD's DC output shows no power. This is really strange and bad.

  1741   Tue Jul 14 00:32:46 2009 rob, albertoOmnistructureEnvironmentRemoval of the cold air deflection device for the MOPA chiller

Quote:

Quote:
Around 2 PM today, I removed the blue flap which has been deflecting the cold air from the AC down into the laser chiller.
Let's watch the laser trends for a few days to see if there's any effect.


Alberto has moved us to stage 2 of this experiment: turning off the AC.

The situation at the control room computers with the AC on minus the blue flap is untenable--it's too cold and the air flow has an unpleasant eye-drying effect.


I turned the AC back on because the temperature of the room was going up so also that of the laser chiller.
  1740   Mon Jul 13 23:03:14 2009 rob, albertoOmnistructureEnvironmentRemoval of the cold air deflection device for the MOPA chiller

Quote:
Around 2 PM today, I removed the blue flap which has been deflecting the cold air from the AC down into the laser chiller.
Let's watch the laser trends for a few days to see if there's any effect.


Alberto has moved us to stage 2 of this experiment: turning off the AC.

The situation at the control room computers with the AC on minus the blue flap is untenable--it's too cold and the air flow has an unpleasant eye-drying effect.
  1739   Mon Jul 13 16:59:10 2009 ZachUpdate GigE Phase Camera

Today, I moved the router from on top of the PSL into the control room in order to perform dark field tests on the GC650 (which I also moved).  The GC750 along with the lens that was on it and the mount it was on has been lent to Ricardo's lab for the time being.  I successfully triggered the GC650 externally and I also characterized the average electronic noise.  For exposure times less than 1 microsecond, the average noise contribution appears to be a constant 15 on a 12-bit scale.

  1738   Mon Jul 13 15:48:05 2009 ranaOmnistructureEnvironmentRemoval of the cold air deflection device for the MOPA chiller
Around 2 PM today, I removed the blue flap which has been deflecting the cold air from the AC down into the laser chiller.
Let's watch the laser trends for a few days to see if there's any effect.
  1737   Mon Jul 13 15:14:57 2009 AlbertoUpdateComputersDAQAWG

Today Alex came over, performed his magic rituals on the DAQAWG computer and fixed it. Now it's up and running again.

I asked him what he did, but he's not sure of what fixed it. He couldn't remember exactly but he said that he poked around, did something somewhere somehow, maybe he tinkered with tpman and eventually the computer went up again.

Now everything is fine.

  1736   Mon Jul 13 00:53:50 2009 AlbertoDAQComputersAll computers down

Quote:

Quote:

I popped by the 40m, and was dismayed to find that all of the front end computers are red (only framebuilder, DAQcontroler, PEMdcu, and c1susvmw1 are green....all the rest are RED).

 

I keyed the crates, and did the telnet.....startup.cmd business on them, and on c1asc I also pushed the little reset button on the physical computer and tried the telnet....startup.cmd stuff again.  Utter failure. 

 

I have to pick someone up from the airport, but I'll be back in an hour or two to see what more I can do.

 I think the problem was caused by a failure of the RFM network: the RFM MEDM screen showed frozen values even when I was power recycling any of the FE computers. So I tried the following things:

- resetting the RFM switch
- power cycling the FE computers
- rebooting the framebuilder
 
but none of them worked.  The FEs didn't come back. Then I reset C1DCU1 and power cycled C1DAQCTRL.
 
After that, I could restart the FEs by power recycling them again. They all came up again except for C1DAQADW. Neither the remote reboot or the power cycling could bring it up.
 
After every attempt of restarting it its lights on the DAQ MEDM  screen turned green only for a fraction of a second and then became red again.
 
So far every attempt to reanimate it failed.

 

After Alberto's bootfest which was more successful than mine, I tried powercycling the AWG crate one more time.  No success.  Just as Alberto had gotten, I got the DAQ screen's AWG lights to flash green, then go back to red.  At Alberto's suggestion, I also gave the physical reset button another try.  Another round of flash-green-back-red ensued.

When I was in a few hours ago while everything was hosed, all the other computer's 'lights' on the DAQ screen were solid red, but the two AWG lights were flashing between green and red, even though I was power cycling the other computers, not touching the AWG at the time.  Those are the lights which are now solid red, except for a quick flash of green right after a reboot.

I poked around in the history of the curren and old elogs, and haven't found anything referring to this crazy blinking between good and bad-ness for the AWG computers.  I don't know if this happens when the tpman goes funky (which is referred to a lot in the annals of the elog in the same entries as the AWG needing rebooting) and no one mentions it, or if this is a new problem.  Alberto and I have decided to get Alex/someone involved in this, because we've exhausted our ideas. 

  1735   Mon Jul 13 00:34:37 2009 AlbertoDAQComputersAll computers down

Quote:

I popped by the 40m, and was dismayed to find that all of the front end computers are red (only framebuilder, DAQcontroler, PEMdcu, and c1susvmw1 are green....all the rest are RED).

 

I keyed the crates, and did the telnet.....startup.cmd business on them, and on c1asc I also pushed the little reset button on the physical computer and tried the telnet....startup.cmd stuff again.  Utter failure. 

 

I have to pick someone up from the airport, but I'll be back in an hour or two to see what more I can do.

 I think the problem was caused by a failure of the RFM network: the RFM MEDM screen showed frozen values even when I was power recycling any of the FE computers. So I tried the following things:

- resetting the RFM switch
- power cycling the FE computers
- rebooting the framebuilder
 
but none of them worked.  The FEs didn't come back. Then I reset C1DCU1 and power cycled C1DAQCTRL.
 
After that, I could restart the FEs by power recycling them again. They all came up again except for C1DAQADW. Neither the remote reboot or the power cycling could bring it up.
 
After every attempt of restarting it its lights on the DAQ MEDM  screen turned green only for a fraction of a second and then became red again.
 
So far every attempt to reanimate it failed.
  1734   Sun Jul 12 23:14:56 2009 JenneOmnistructureGeneralWeb screenshots aren't being updated

Before heading back to the 40m to check on the computer situation, I thought I'd check the web screenshots page that Kakeru worked on, and it looks like none of the screens have been updated since June 1st.  I don't know what the story is on that one, or how to fix it, but it'd be handy if it were fixed.

  1733   Sun Jul 12 20:06:44 2009 JenneDAQComputersAll computers down

I popped by the 40m, and was dismayed to find that all of the front end computers are red (only framebuilder, DAQcontroler, PEMdcu, and c1susvmw1 are green....all the rest are RED).

 

I keyed the crates, and did the telnet.....startup.cmd business on them, and on c1asc I also pushed the little reset button on the physical computer and tried the telnet....startup.cmd stuff again.  Utter failure. 

 

I have to pick someone up from the airport, but I'll be back in an hour or two to see what more I can do.

  1732   Sun Jul 12 20:05:06 2009 ranaOmnistructureEnvironmentChanged office temp
This is a 7-day minute trend. There's no obvious effect in any of the channels looking back 2 days.

Seems like the laser chiller fix has made the laser much more immune to the office area temperature.
Attachment 1: Picture_3.png
Picture_3.png
  1731   Fri Jul 10 19:56:23 2009 rana, kojiOmnistructureEnvironmentChanged office temp

I have increased the temperature setpoint in the office area by ~0.75 deg F. Figure attached. Also a few days ago I increased the setpoint of the AC in the control room. Looks like the Laser is able to handle the changes in office area temperature so far, but lets see how it fares over the weekend.

Attachment 1: therm.JPG
therm.JPG
  1730   Fri Jul 10 17:32:08 2009 ranaUpdateEnvironmentseisBLRMS & mafalda restarted
Rana, Alberto

Mafalda's ethernet cable had fallen out of the connector on the hub-side. We reconnected it and rebooted mafalda and restarted seisBLRMS.

Mafalda didn't mount /cvs/cds on start up for some reason. I mounted it using 'sudo mount linux1:/home/cds /cvs/cds' and it took at least
30 seconds, so maybe there's a timeout issue. Seems OK now.
  1729   Thu Jul 9 19:24:50 2009 ranaUpdatePEMmore mic position changes; mics not picking up high frequencies
Might be the insidious 850 Hz AA filters in the black AA box which precedes the ADC.

Dan Busby fixed up the PSL/IOO chassis. WE might need to do the same for the PEM stuff.
  1728   Thu Jul 9 19:05:32 2009 ClaraUpdatePEMmore mic position changes; mics not picking up high frequencies

Bonnie has been strung up on bungees in the PSL so that her position/orientation can be adjusted however we like. She is now hanging pretty low over the table, rather than being attached to the hanging equipment shelf thing. Butch Cassidy has been hung over the AS table.

Moving Bonnie increased the coherence for the PMC_ERR_F signal, but not the MC_L. Butch Cassidy doesn't have much coherence with either.

I noticed that the coherence would drop off very sharply just after 10 kHz - there would be no further spikes or anything of the sort. I used my computer to play a swept sine wave (sweeping from 20Hz to 10kHz) next to Butch Cassidy to see if the same drop-off occurred in the microphone signal itself. Sure enough, the power spectrum showed a sharp drop around 10kHz. Thinking that the issue was that the voltage dividers had too high impedance, I remade one of them with two 280 Ohm and one 10 Ohm resistor, but that didn't make any difference. So, I'm not sure what's happening exactly. I didn't redo the other voltage divider, so Bonnie is currently not operating.

 

Attachment 1: DSC_0569.JPG
DSC_0569.JPG
Attachment 2: DSC_0570.JPG
DSC_0570.JPG
Attachment 3: bonnie_psl_hi_mcl12.pdf
bonnie_psl_hi_mcl12.pdf
Attachment 4: bonnie_psl_hi_errf12.pdf
bonnie_psl_hi_errf12.pdf
Attachment 5: bc_as_table.pdf
bc_as_table.pdf
Attachment 6: powerspec.pdf
powerspec.pdf
  1727   Thu Jul 9 02:18:09 2009 ranaUpdatePEMBonnie moved to PSL, getting some coherence with the PMC_ERR channel
My guess is that we need a different acoustic strategy. The microphones are mainly for high frequencies,
so you should not expect any coherence with MC_L (or even better, MC_F) below 100 Hz. I expect that the
main coherence for MC_F will come from the PSL in that band.

After subtracting that one out, we should look at the signal from the lock of the X or Y arm, and see
if we can nail that by putting a mic right above the AS table (leaving enough room to take the lid off).
If that works OK, we can find a spot under the lid and se if it gets better.
  1726   Wed Jul 8 19:42:37 2009 ClaraUpdatePEMBonnie moved to PSL, getting some coherence with the PMC_ERR channel

In her position overlooking whichever table it is that is next to the PSL, Bonnie drummed up some decent coherence with the PSL-PMC_ERR channel, but not so much with the MC_L. I moved her into the PSL itself, and now there is rather good coherence with the PMC_ERR channel, but still not so great for MC_L.

DSC_0567.JPG

Bonnie's new home in the PSL.

Attachment 2: bonnie2_pmcerr.pdf
bonnie2_pmcerr.pdf
Attachment 3: bonnie_PSL.pdf
bonnie_PSL.pdf
  1725   Wed Jul 8 19:13:19 2009 AlbertoUpdatePSLPSL beam aligned to the Mode Cleaner

Today I tuned the periscope on the PSL table to align the beam to the Mode Cleaner. With the Wave Front Sensor control off, I minimized the reflection from the MC and maximized the transmission. While doing that I also checked that the transmitted beam after the MC didn't lose the alignment with the interferometer's main Faraday isolator.

In this way, I've got a reflection, as read from the MC_REFLPD_MC, of about 0.6. Then I centered the WFS on the AS table. After that the WFS alignment control brought the reflection to 0.25 and a nice centered bull-eye spot showed on the monitor.

  1724   Wed Jul 8 18:46:56 2009 DmassAoGElectronicsBeam Scan Funky

The beam scan (which has been living in the bridge subbasement for a bit now) is in a state of imperfection.

I noticed that:

  • The waist reading seems to change by not insignificant amounts as you move the spot across the head, even for just small perturbations about the center.
  • None of the features which require two slits seem to be working (unsure if this is software or hardware related)

I took some pictures to try and illuminate the situation - The inverted images are included to make it easier to see the flecks (?) in the slits

I am not sure how to figure out if any bit of the scan is/has been fried.

 

Pending further investigation, enjoy large error bars in your scan measurements!

 

PICTURES OF BOTH SLITS ON THE BEAMSCAN HEAD:

Attachment 1: beamscanhead3.png
beamscanhead3.png
Attachment 2: beamscanhead6.png
beamscanhead6.png
  1723   Wed Jul 8 12:36:15 2009 ClaraUpdatePEMCoherences and things

After setting up the microphones last week, I modified the Wiener filtering programs so as to include the microphone signals. They didn't seem to do much of anything to reduce the MC_L signal, so I looked at coherences. The microphones don't seem to have much coherence with the MC_L signal at all. I tried moving Bonnie to near the optical table next to the PSL (which isn't in a vacuum, and thus would, presumably, be more affected by acoustic noise), but that didn't seem to make much of a difference. Eventually, I'd like to put a mic in the PSL itself, but I need to work out how to mount it first.

DSC_0564.JPG

Bonnie's new location.

You can see in bonnie_butch.pdf that none of the mic signals are giving very good coherence, although they all seem to have a peak at 24 Hz. (In fact, everything seems to have a peak there. Must be a resonant frequency of something in the mode cleaner.)

I've also attached plots of the coherences for all six accelerometers and the three Guralp seismometer axes. I plotted the most coherent traces together in the last pdf: the y-axes of the MC2 accelerometer and the two seismometers (the Ranger measures ONLY y) and, interestingly, the z-axis of the MC2 accelerometer. Unsurprisingly, the seismometers are most coherent at the low frequencies, and the MC2-Y accelerometer seems to be coherent at very similar frequencies. The MC2-Z accelerometer, on the other hand, seems to be coherent at the higher frequencies, and is highly complementary to the others. I am not really sure why this would be...

Finally, I was curious about how the noise varies throughout the day, because I didn't want to mistakenly decide that some particular configuration of accelerometers/seismometers/whatever was better than another b/c I picked the wrong time of day to collect the data. So, here is a plot of Wiener filters (using only accelerometer data) taken over 2-hour intervals throughout the entirety of July 6, 2009 (midnight-midnight local).

2hr_allday_1.png

It's a little bit confusing, and I should probably try to select some representative curves and eliminate the rest to simplify things, but I don't have time to do that before the meeting, so this will have to suffice for now.

Attachment 2: bonnie_butch.pdf
bonnie_butch.pdf
Attachment 3: Gseis_100.pdf
Gseis_100.pdf
Attachment 4: mc1_xyz.pdf
mc1_xyz.pdf
Attachment 5: mc2_xyz.pdf
mc2_xyz.pdf
Attachment 6: most_coherent.pdf
most_coherent.pdf
  1722   Wed Jul 8 11:13:36 2009 AlbertoOmnistructureComputerswireless router disconnected

Once again, this morning I found the wireless router disconnected from the LAN cable. No martian WiFi was available.

I wonder who is been doing that and for what reason.

  1721   Wed Jul 8 11:08:43 2009 ZachUpdateCamerasGigE Phase Camera

The plan for the optical setup has been corrected after it was realized that it would be impossible to isolate a 29.501 MHz frequency from a 29.499 MHz one because they are so close in value.  Instead, we decided to adopt the setup pictured below.  In this way, the low-pass filter should have no trouble isolating 29.501-29.5 MHz from 29.501 + 29.5 MHz.  Also, we decided to scrap the idea of sending Alberto's laser through a fiber optic cable after hearing rumors of extra lasers.  Since I shouldn't have to share a beam when the second laser comes in, I plan on setting up both lasers on the same optics bench.  I've been working on the software while waiting for supplies, but I should be able to start building the trigger box today (assuming the four-pair cable is delivered).

Attachment 1: fig1.pdf
fig1.pdf
Attachment 2: fig2.pdf
fig2.pdf
  1720   Wed Jul 8 11:05:40 2009 Chris ZimmermanUpdateGeneralWeek 3/4 Update

The last week I've spent mostly working on calculating shot noise and other sensitivities in three michelson sensor setups, the standard michelson, the "long range" michelson (with wave plates), and the proposed EUCLID setup.  The goal is to show that there is some inherent advantage to the latter two setups as displacement sensors.  This involved looking into polarization and optics a lot more, so I've been spending a lot of time on that also.  For example, the displacement sensitivity/shot noise on the standard michelson is around 6:805*10^-17 m/rHz at L_=1*10^-7m, as shown in the graph.  NSD_Displacement.png

  1719   Wed Jul 8 10:56:04 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

This week, I've been working on adapting the last week's circuit to make it buildable. Mostly this has involved picking components that are already in the lab, adding tunable components when necessary, and planning roughly how the components should be laid out on a board. I then built the circuit and put it in a box with BNC connectors for easy connection during testing. A picture of the built circuit is attached.

For initial testing, the transformer was removed from the design; since this changed the response of the circuit, I added two resistors to correct the response. A figure showing a schematic of the built circuit is attached. The expected responce of the circuit is also shown; the magnitude (solid) and phase (dashed) of the voltage across the EOM are shown in green, and the impedance of the circuit is shown in blue. While this response has sharp peaks and 50 Ohms (34 dB) of impedance at resonances, the gain is low compared to the circuit with the transformer. This means that, as is, this circuit cannot be used to drive the EOM; it is simply for testing purposes.

Attachment 1: DSC_0566.JPG
DSC_0566.JPG
Attachment 2: InitialBuiltCkt.pdf
InitialBuiltCkt.pdf
Attachment 3: BuiltCkt_ExpectedResponse.png
BuiltCkt_ExpectedResponse.png
  1718   Tue Jul 7 16:06:59 2009 ClaraUpdateComputer Scripts / ProgramsDTT synchronization errors, help would be appreciated

I am attempting to use the DTT program to look at the coherence of the individual accelerometer signals with the MC_L signal. Rana suggested that I might break up the XYZ configuration, so i wanted to see how the coherence changed when I moved things around over the past couple of weeks, but I keep getting a synchronization error every time I try to set the start time to more than about 3 days ago. I tried restarting the program and checking the "reconnect" option in the "Input" tab, neither of which made any kind of difference. I can access this data with no problem from the Data Viewer and the Matlab scripts, so I'm not really sure what is happening. Help?

EDIT: Problem solved - Full data was not stored for the time I needed to access it for DTT.

  1717   Tue Jul 7 15:08:49 2009 KojiSummaryPhotos40 high school students visited 40M

Alan and Alberto conducted a tour of 40 high-school students.
It may be the same tour that Rana found a spare PMC during the tour explanation as far as I remember...

Attachment 1: IMG_1848.jpg
IMG_1848.jpg
  1716   Thu Jul 2 17:37:36 2009 steveUpdateVACV1 interlock test

Joe, Alberto and Steve

We tested gate valve V1 interlock by :

1, decelerated rotation by brake from maglev controller unit.

2, turned maglev  controller off from controller unit.

3, unpluged 220VAC plug from wall socket

None of the above action triggered V1 to close. This needs to be corrected in the future.

The MEDM monitor screen of maglev indicated the correct condition changes.

 

 

  1715   Thu Jul 2 16:45:06 2009 ClaraUpdatePEMBonnie and Clyde are officially in operation! (Butch Cassidy and the Sundance Kid are in temp position)

I hooked up Bonnie and Clyde last night and tested it today. First I tried some loud noises to make sure I could identify them on the readout. Then, Steve suggested I try to look for some periodic stuff. I set up Butch Cassidy and the Sundance Kid on the cabinets by the MC2 optic. Now for graphs!

 

bonnie_test_marked.png

I tapped on the microphone a few times. I also yelled a bit, but this is sampling by seconds, so perhaps they got overwhelmed by the tapping.

bonnie_test2_marked.png

This time I tried some more isolated yells. I started with a tap so I'd be sure to be able to recognize what happened. Apparently, not so necessary.

bonniebutch_2sbeat_marked.png

Here, it looks like a pretty strong periodic pattern on the second mic (Butch Cassidy). I replaced the lines with dashed ones where the pattern was a little less clear. Possibility interference from something. Mic1 (Bonnie) seems to show a pretty regular beat pattern, which seems reasonable, as it isn't particularly close to any one instrument fan.

 

So, anyway. I thought those were neat. And that I wanted to share.

 

  1714   Thu Jul 2 06:31:35 2009 ClaraUpdatePEMFirst mic in place and connected

I clamped Bonnie (microphone) to the top of a chamber near the vertex of the arms and placed Clyde (pre-amp) on the table right below (see picture). The cable was laid and Bonnie and Clyde are plugged into port #13 on the ADC. The second cable was plugged into port #14, but it is not connected to anything. I placed the looped up cable on top of the cabinet holding the ADC.

Note: the angle in the photograph is such that we are looking along the y-arm.


Attachment 1: bonnieandclyde.JPG
bonnieandclyde.JPG
  1713   Thu Jul 2 05:27:12 2009 ClaraUpdatePEMVoltage Divider Oops

I tested the voltage dividers and was getting up to about 3V. I retested the mic w/o the voltage divider in place, and, lo and behold, I was able to generate about 70-75V (previously, I maxed out at 45V). I'm not 100% sure why this was, but it occurs to me that, before, the sounds I was generating were short in duration (loud claps, short yelps). This time, I tried yelling continuously into the microphone. So, probably, I simply wasn't seeing the real peak before on the scope because it was too short to pick up. I have corrected the voltage dividers (by replacing the 25.5 kOhm resistors, which were in parallel with the ADC, with 10 kOhm resistors, taking the voltage ratio to ~60:1) and tested them. I haven't been able to generate more than 1500 mV, so I think they are safe. (It's possible we would have been fine with the old setup, since I think it would be hard to get any noises as loud as I was making, but better safe than sorry, right?)

I'm attaching a diagram of the new-and-improved voltage dividers.

voltage_divider_diagram.png

  1712   Wed Jul 1 11:04:27 2009 ZachUpdateCamerasGigE Phase Camera

This past week, I have building a sine wave rectifier and trying to write a simple program that displays a ccd image to familiarize myself with the code.  I also wrote a progress report in which I included the following images of the sine wave rectifier circuit as well as the optical chain including the phase-locked loop.  The hirose connector arrived so I can begin soldering the electronics together and testing the trigger box with the ccd.  I am waiting on the universal PDH box as well as another fiber coupler to begin setting up the optics.  In order to avoid the frustrations associated with sending a laser beam down a long pipe to an optical bench across the room, I will be transmitting laser 1 to the ccd by means of a fiber optic cable and dealing with the alternative new and exciting frustrations.

 

Attachment 1: trigger.jpg
trigger.jpg
Attachment 2: fig1b.pdf
fig1b.pdf
  1711   Wed Jul 1 11:00:52 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

Since last week, I have come up with a new circuit, which is shown in the attached figure. The magnitude (solid) and phase (dashed) of the voltage across the EOM (red), the ratio between the voltage across the EOM and the voltage across the primary nodes of the transformer (blue), and the impedance through the primary port of the transformer (red) are also shown in an attached figure. As can be seen on the plot, resonance occurs at 11 MHz, 29.5 MHz, and 55 MHz, as specified. Also, at these resonant frequencies, the impedance is about 50 Ohms (34 dB). The gain between the voltage across the EOM and the voltage across the primary nodes of the transformer (or output of the crystal oscillator) is about 12 dB; we'd like a higher gain than this, but this gain is primarily governed by the ratio between the secondary and primary inductances in the transformer, and we are using the largest available ratio (on the Coilcraft website) that has the necessary bandwidth. Because of this, we will likely have to add another component between the crystal oscillator and the EOM circuit, to get the voltage to the desired 8.5 Vp across the EOM (for an optical modulation depth of 0.1 rad).

The current and power through the primary port of the tranformer are 43-85 mA and 25-92 mW, respectively. Since the transformer ratings are 250 mA and 1/4 W for current and power; these values should be safe to use with the intended transformer. Also, the highest power dissipated by a resistor in the circuit (not including the 50 Ohm resistor that is part of the crystal oscillator setup) is around 74 mW.

Attachment 1: EOMCKT.png
EOMCKT.png
Attachment 2: PR1_VoltagePlot.pdf
PR1_VoltagePlot.pdf
  1710   Wed Jul 1 10:56:42 2009 Chris ZimmermanUpdateGeneralWeek 2/3 Update

I spent the last week working a lot with the differences between a basic Michelson readout and the new one as a displacement sensor.  The new one (w/ wave plates) ends with two differently polarized beams and should have better sensitivity; I've also been going through noise/sensitivity calculations for each, although that hit a road block when I had to start the 1st SURF progress report, which has taken up most of my time since Saturday.

  1709   Tue Jun 30 23:09:40 2009 AlbertoUpdateLockingchronicles of some locking attempts

Tonight I tried to lock the interferometer. At the first attempts the arm power didn't go above about 4. The mode cleaner seemed to be not well aligned and it lost lock or got stuck on a wrong mode. I had to run the MC_UP and MC_DOWN scripts to lock it again.

After that the locking proceed more smoothly; at least till a power level in the arms of about 60. Then again the mode cleaner lost lock and I had to run the scripts again. Without the MCWFS servo off the MC reflected power is still rather high (about 1.7); also even when the WFS servo is engaged the reflected power is about 0.5, versus 0.3 that it should be.

Those are both signs of a not very good alignment. Tomorrow I'll have to work on the injection periscope on the PSL table to try to fix that.

  1708   Sat Jun 27 03:16:16 2009 ClaraUpdatePEMExciting microphone things!

So, I'm double-posting, but I figured the last post was long enough as it was, and this is about something different. After double and triple checking the XLR cables, I hooked up the microphone setup (mic---preamp---output) to the oscilloscope to figure out what kind of voltage would register with loud noises. So, I clapped and shouted and forgot to warn the other people in the lab what I was doing (sorry guys) and discovered that, even on the lowest gain setting, my loud noises were generation 2-3 times as much voltage as the ADC can handle (2V). And, since our XLR cables are so freaking long, we probably want to go for a higher gain, which puts us at something like 20 times too much voltage. I doubt this is really necessary, but it's late (early) and I got camera-happy, so I'm going to share anyway:

mic_voltage_out.png

So, to deal with this issue, I made some nifty voltage dividers. Hopefully they are small enough to fit side-by-side in the ports without needing extra cableage. Anyway, they should prevent the voltage from getting larger than 2V at the output even if the mic setup is producing 50V. Seeing as my screaming as loud as I could about 2mm away from the mic at full gain could only produce 45V, I think this should be pretty safe. I put the ADC in parallel with a 25.5 kOhm resistor, which should have a noise like 10^-8 V/rHz. This is a lot smaller than 1 uV/rHz (the noise in the ADC, if I understood Rana's explanation correctly), so the voltage dividers should pose a noise issue. Now for pictures.

voltage_dividers.png

I opened one so you can see its innards.

voltage_divider_sketch.png

In case the diagram on the box was too small to decipher...

And finally, I came up with a name scheme for the mics and pre-amps. We now have two Bluebird (bacteriophage) mics named Bonnie and Butch Cassidy. Their preamps are, naturally, Clyde and The Sundance Kid. Sadly, no photos. I know it's disappointing. Also, before anyone gives me crap for putting the labels on the mics upside-down, they are meant to be hung or mounted from high things, and the location (and stiffness) of the cable prevents us from simply standing them up. So they will more than likely be in some kind of upsidedownish position.

  1707   Sat Jun 27 02:48:09 2009 ClaraUpdatePEMI moved accelerometers and made some pretty pictures

I have been working on finding the best spots to put the accelerometer sets in order to best subtract out noise (seismometers next!). Here is a plot of what I've done so far:

80min_accel_0123.png

All of these were 80-minute samples. The dashed line is unfiltered, solid line filtered. So, Setup #1 looks the best so far, but I didn't leave it there very long, so perhaps it was just a really awesome 80 minutes. I've put the accelerometers back in the Setup #1 position to make sure that it is really better.

And, in case you can't intuitively figure out what configuration the accelerometers are in by such descriptive names, here are some helpful pictures. I didn't know about the digital cameras at first, so these are actually sketches from my notebook, which I helpfully labeled with the setup numbers, color-coded to match the graph above! Also, there are some real-life photographs of the current arrangement (Setup #1' if you forgot).

MC1_accel_sketch_side.png

MC1_accel_sketch_top.png

MC2_accel_sketch.png

Doesn't this one look kind of Quentin Blake-esque? (He illustrated for Roald Dahl.)

MC1acc_S01.JPG

This is the MC1 set.

MC2acc_S123.JPG

Guess which one this is!

  1706   Fri Jun 26 19:14:04 2009 ranaUpdateEnvironmentseisBLRMS for the past 3 weeks
Restarted the seisBLRMS.m on mafalda (running a term on op540m). Don't know why it stopped - the
terminal had a 'disabled by EPICS' message even though the EPICS enable button was enabled.

I also changed the delay from 4 to 2 minutes. So now it is calculating a 64 s PSD starting from 2 minutes ago. We had
put it back to 4 minutes long ago since the framebuilder was acting slow, but maybe it will work now since its using
the NDS1 protocol instead of direct frame reading. If it starts not finding data, please kill the seisBLRMS and restart
it with a 3 minute delay.
  1705   Fri Jun 26 18:00:36 2009 JenneUpdatePEMNew PEM channels for the fancy-pants new microphones

[ Jenne, Clara ]

We made new channels for the microphones which came in this week, by editing C1ADCU_PEM.ini (and making an appropriate backup before we modified it) then restarting the framebuilder and the frontend computer C0DCU1.  The new channels are:

C1:PEM-AUDIO_MIC1

C1:PEM-AUDIO_MIC2

These are connected to channels 13 and 14 on the PEM ADCU board, just next to the GURALP seismometer channels.

 

Clara is testing the mics so the max output voltage can be limited to +-2V for the DAQ, then we'll hook them up to our new channels and listen to the IFO (and all the audio frequency noises around it).

  1704   Fri Jun 26 15:22:28 2009 ClaraUpdatePEMXLR cables tested and labeled

We now have two 80-foot, female-to-female XLR cables for our pretty new microphones, one yellow and one purple. They have been tested and appropriately labeled.

Also, here is a very helpful pdf for how to properly attach the XLR connectors to a raw quad cable, as well as one for how to put the actual connectors together (ignore the cable instructions on the connector page... the cable depicted is not a quad cable).

 

 

Attachment 1: NC3FXX-EMC.pdf
NC3FXX-EMC.pdf
Attachment 2: Cat11_p35.pdf
Cat11_p35.pdf
  1703   Thu Jun 25 21:00:30 2009 ClaraUpdatePEMMC1 Accelerator set moved again; new XLR cables

I moved the MC1 set of accelerators. Might have bumped things. If things aren't working, look around the MC1 chamber.

Also, I constructed two new XLR cables, but have not tested them yet.

  1702   Thu Jun 25 17:27:42 2009 ranaUpdateComputerstdsresp on linux
I downloaded the tdsresp.pl from LLO and put it into the various TDS/bin paths. Also updated the LLO specific path stuff. It runs.
  1701   Thu Jun 25 10:28:58 2009 steveHowToVACCryopump is regenerated
The Cryopump's VC1 valve to IFO was closed yesterday.
The compressor  Helium pressure was 235 PSI. The cold head temp on H2 vapor pressure gauge was reading ~14 Kelvin,
The compressor and piston driver were turned off to let cold head warm up to room temp.
The flow path from Cryo to TP3 were checked to insure that only VC2 and V5 would be open for pumping.
VC2 valve was opened to TP3 through V5
Now as the Cryo was warming up while TP3 drag turbo pump was pumping away the accumulated ice, that was melting and vaporizing.
This is shown on one day the plot below.
To check outgassing rate of the Cryo pump after one day of pumping V5 was closed for 20 minutes.
The accumulation was 1.3 mTorr in 20 min
This means the Cryo is clean, it is ready to be started up in the future.
VC2 was closed to seal this condition.
The flow path between VC2, VM3, V7, V6 , VA6 and manual needle valve  would be pumped for one day through V5 to TP3 to clean up 
Attachment 1: Creg.jpg
Creg.jpg
  1700   Wed Jun 24 17:08:16 2009 steveBureaucracySAFETYsurf student safety

The 40m lab specific safety training is done. The participants were

Stephanie Erickson, Clara Bennett, Chris Zimmerman, Zach Commings, Michelle Stephen surfs and Drew Cappel  postock.

They have already went through the Caltech Safety Office laser and general safety training.

They still have to read, understand and sign the the SOP for the laser & lab

  1699   Wed Jun 24 14:43:19 2009 robUpdateComputerstdsresp on linux => pzresp

 

tdsresp is broken on our linux control room machines.  I made a little perl replacement which uses the DiagTools.pm perl module, called pzresp.  It's in the $SCRIPTS/general directory, and so in the path of all the machines.  I also edited the cshrc.40m file so that on linux machines tdsresp points to this perl replacement.

I've patched DiagTools.pm to circumnavigate the tdsdmd bug described here.  I also added a function to DiagTools.pm called diagRespNoLog, which is just like diagResp but without that pesky log file.

 


Here's the output from the tdsresp binary on CentOS:
allegra:~>tdsresp 941.54 10000 100 10 C1:LSC-ITMX_EXC C1:LSC-PD1_Q C1:LSC-PD1_I
nan nan nan nan nan nan nan
nan nan nan nan nan nan nan
nan nan nan nan nan nan nan
*** glibc detected *** tdsresp: free(): invalid next size (fast): 0x089483e8 ***
======= Backtrace: =========
/lib/libc.so.6[0xa800f1]
/lib/libc.so.6(cfree+0x90)[0xa83bc0]
/usr/lib/libstdc++.so.6(_ZdlPv+0x21)[0xf7f36571]
tdsresp[0x8057fbb]
tdsresp[0x805b394]
/lib/libc.so.6(__libc_start_main+0xdc)[0xa2ce8c]
tdsresp(__gxx_personality_v0+0x169)[0x804ddd1]
======= Memory map: ========
00242000-00249000 r-xp 00000000 fd:00 15400987                           /lib/librt-2.5.so
00249000-0024a000 r--p 00006000 fd:00 15400987                           /lib/librt-2.5.so
0024a000-0024b000 rw-p 00007000 fd:00 15400987                           /lib/librt-2.5.so
009f9000-00a13000 r-xp 00000000 fd:00 15400963                           /lib/ld-2.5.so
00a13000-00a14000 r--p 00019000 fd:00 15400963                           /lib/ld-2.5.so
00a14000-00a15000 rw-p 0001a000 fd:00 15400963                           /lib/ld-2.5.so
00a17000-00b55000 r-xp 00000000 fd:00 15400974                           /lib/libc-2.5.so
00b55000-00b57000 r--p 0013e000 fd:00 15400974                           /lib/libc-2.5.so
00b57000-00b58000 rw-p 00140000 fd:00 15400974                           /lib/libc-2.5.so
00b58000-00b5b000 rw-p 00b58000 00:00 0 
00b5d000-00b70000 r-xp 00000000 fd:00 15400984                           /lib/libpthread-2.5.so
00b70000-00b71000 r--p 00012000 fd:00 15400984                           /lib/libpthread-2.5.so
00b71000-00b72000 rw-p 00013000 fd:00 15400984                           /lib/libpthread-2.5.so
00b72000-00b74000 rw-p 00b72000 00:00 0 
00b76000-00b78000 r-xp 00000000 fd:00 15400981                           /lib/libdl-2.5.so
00b78000-00b79000 r--p 00001000 fd:00 15400981                           /lib/libdl-2.5.so
00b79000-00b7a000 rw-p 00002000 fd:00 15400981                           /lib/libdl-2.5.so
00b7c000-00ba1000 r-xp 00000000 fd:00 15400975                           /lib/libm-2.5.so
00ba1000-00ba2000 r--p 00024000 fd:00 15400975                           /lib/libm-2.5.so
00ba2000-00ba3000 rw-p 00025000 fd:00 15400975                           /lib/libm-2.5.so
00bca000-00bdd000 r-xp 00000000 fd:00 15401011                           /lib/libnsl-2.5.so
00bdd000-00bde000 r--p 00012000 fd:00 15401011                           /lib/libnsl-2.5.so
00bde000-00bdf000 rw-p 00013000 fd:00 15401011                           /lib/libnsl-2.5.so
00bdf000-00be1000 rw-p 00bdf000 00:00 0 
00dca000-00dd5000 r-xp 00000000 fd:00 15400986                           /lib/libgcc_s-4.1.2-20080825.so.1
00dd5000-00dd6000 rw-p 0000a000 fd:00 15400986                           /lib/libgcc_s-4.1.2-20080825.so.1
08048000-080b7000 r-xp 00000000 00:17 6455328                            /cvs/cds/caltech/apps/linux/tds/bin/tdsresp
080b7000-080ba000 rw-p 0006e000 00:17 6455328                            /cvs/cds/caltech/apps/linux/tds/bin/tdsresp
080ba000-080bb000 rw-p 080ba000 00:00 0 
0893d000-0896b000 rw-p 0893d000 00:00 0                                  [heap]
f5e73000-f5e74000 ---p f5e73000 00:00 0 
f5e74000-f6874000 rw-p f5e74000 00:00 0 
f692d000-f6931000 r-xp 00000000 fd:00 15400995                           /lib/libnss_dns-2.5.so
f6931000-f6932000 r--p 00003000 fd:00 15400995                           /lib/libnss_dns-2.5.so
f6932000-f6933000 rw-p 00004000 fd:00 15400995                           /lib/libnss_dns-2.5.so
f6956000-f6a12000 rw-p f6a31000 00:00 0 
f6a74000-f6a7d000 r-xp 00000000 fd:00 15400997                           /lib/libnss_files-2.5.so
f6a7d000-f6a7e000 r--p 00008000 fd:00 15400997                           /lib/libnss_files-2.5.so
f6a7e000-f6a7f000 rw-p 00009000 fd:00 15400997                           /lib/libnss_files-2.5.so
f6a7f000-f6a80000 ---p f6a7f000 00:00 0 
f6a80000-f7480000 rw-p f6a80000 00:00 0 
f7480000-f7481000 ---p f7480000 00:00 0 
f7481000-f7e83000 rw-p f7481000 00:00 0 
f7e83000-f7f63000 r-xp 00000000 fd:00 6236924                            /usr/lib/libstdc++.so.6.0.8
f7f63000-f7f67000 r--p 000df000 fd:00 6236924                            /usr/libAbort

 
  1698   Wed Jun 24 12:09:24 2009 ClaraUpdatePEMWeek 1(ish)

I spent the week reading up on filter algorithm theory, particularly Wiener filtering. I have also learned how to get data from specific channels at specific times, and I've been getting myself acquainted with Matlab (which I have not previously used). Finally, I started messing around with the positioning of the accelerometers and seismometers in order to try to find the setup that yields the best filtration.

  1697   Wed Jun 24 12:04:22 2009 ZachUpdateCamerasSURF entry

This week, I've been reading some literature concerning PLL and familiarizing myself with LINUX, MATLAB, and high-pass filter circuits.  In MATLAB, I started constructing matrices to be used for a beam path analysis from the laser output to the ccd camera.  I also built a simple high-pass filter on a bread-board that Joe and I are currently testing with the spectrum analyzer.

  1696   Wed Jun 24 12:04:00 2009 ClaraUpdatePEMaccelerometer clarification

When I said "MC1/MC2 accelerometers," I meant the entire three-axis accelerometer set at each point.

  1695   Wed Jun 24 11:20:40 2009 StephanieUpdateGeneralMultiply Resonant EOM Update

I have created the attached EOM circuit with resonances at 11 MHz, 29.5 MHz, and 55 MHz (the magnitude and phase of the voltage across the EOM are shown in the attached plot). The gain is roughly the same for each resonant peak. Although I have managed to get the impedances at all of the resonant frequencies to equal each other, I am having more trouble getting the impedances to be 50 Ohms (they are currently all around 0.66 Ohms).

For the current circuit, initial calculations show that we will need around 4.7 - 14.2 A of current to drive the EOM at the desired voltage (8 - 24 V); this is much higher than the current rating of most of the available transformers (250 mA), but the necessary current will change as the impedance of the circuit is corrected, so this is probably not a cause for concern. For example, the necessary driving voltages for the current circuit are (2.8 - 8.5 V); if we assume that the 50-Ohm impedance will be purely resistive, then we get a current range of 56 - 170 mA.

Attachment 1: EOM_CktDiagram.JPG
EOM_CktDiagram.JPG
Attachment 2: EOM_VoltagePlot.JPG
EOM_VoltagePlot.JPG
  1694   Wed Jun 24 10:53:34 2009 Chris ZimmermanUpdateGeneralWeek 1/2 Update

I've spent most of the last week doing background reading; fourier transforms, shm, e&m, and other physics that I didn't cover at school.  I also read a few chapters in Saulson, especially the chapter on noise and shot noise.  To get a better grip on what I'm going to be doing I read through the polarization chapter in Hobbs' "Optics" text, mostly on wave plates since that's a large part of this readout.  Since then I've been working up to calculating the shot noise, starting with the electric field throughout the new interferometer readout.

ELOG V3.1.3-